OpenStudy (anonymous):

i got this equation x^2-y^2+z^2-4x-2y-2z+4=0, i reduced it to (x-2)^2-(y+1)^2+(z-1)^2=0. Wolfram told me this is an infinite cone, how do i go about graphing it?

OpenStudy (anonymous):

No idea, but I'd like to bookmark the question.

OpenStudy (anonymous):

you know what, i think i know how to do it. This is an infinite cone, but its shifted, meaning its not on the origin.

OpenStudy (anonymous):

so, i suppose i could set z=0

OpenStudy (anonymous):

then obtain the lines, that would form the cone itself

OpenStudy (jamesj):

What does this cone look like: x^2 + z^2 - y^2 = 0 ? It's a cone that expands along the y-axis and has its apex at the origin.

OpenStudy (anonymous):

right, no problem

OpenStudy (jamesj):

Similarly, your cone expands parallel to the y-axis and has its apex at (2,-1, 1)

OpenStudy (anonymous):

in this case the cone would open on the y-axis cause of the negative sign

OpenStudy (anonymous):

very good

OpenStudy (jamesj):

Yes

OpenStudy (jamesj):

Notice also by the way that the cone is also 'double-sided'; it open up in both directions from the apex.

OpenStudy (anonymous):

so, in reality, there is no need to find the traces, it would appear to be enough to draw intersecting lines through the apex

OpenStudy (anonymous):

i get a circle , with x=0;yz-plane also another circle with y=0;xz-plane then with z=0, there is no graph

OpenStudy (jamesj):

All the cross sections in the xz-plane--for y = constant--will be circles. Stick for the moment with the case x^2 + z^2 - y^2 = 0 For EVERY value of y except y = 0, the locus of the xz points in the xz plane will be circle because the equation above is equivalent to \[ x^2 + z^2 = y^2 > 0 \] Hence if y is a non-zero constant, then the locus of the points (x,y,z) which satisfy that equation is a circle in the xz plane.